How do I solve and show the work for x^2+x-12=0?

Answer by TakeATuition.com(57)

SOLUTION -
X^2 + X - 12 = 0
X^2 + 4X - 3X - 12 = 0

X(X + 4) - 3(X + 4) = 0

(X - 3)(X + 4) = 0
X - 3 = 0

X = 3
X + 4 = 0
X = -4